OpenStudy (anonymous):

Solve x^2 + 9x = –3. (using quadratic formula)

OpenStudy (comm.dan):

Do you know the Quadratic Formula?

OpenStudy (anonymous):

Rearrange to get it in the form \(ax^2 +bx+c=0\) (shouldn't be too hard) and then use \[x = \frac{-b \pm\sqrt{b^2-4ac}}{2a}\] substituting the values that you find. Hope that helps!

OpenStudy (comm.dan):

What is the a, b, and c in this equation? @Britt101

OpenStudy (anonymous):

to start of the function would be \[x^2+9x+3=0\]

OpenStudy (anonymous):

a=1 b=9 c=3

OpenStudy (comm.dan):

Good, now plug in the a, b, and c into the quadratic formula. @Britt101

OpenStudy (comm.dan):

Tell us what you get as your final answer or ask any questions if you have any, as needed. @Britt101

OpenStudy (ryan123345):

Go Luis!

OpenStudy (anonymous):

-9+or-(squared) 69/2?

OpenStudy (anonymous):

@Comm.Dan

OpenStudy (anonymous):

Just if anyone wants to copy this code: [x=\frac{-b \pm \sqrt{b^2-4ac}} {2a}\] Smash in a \ at the beginning

OpenStudy (comm.dan):

I think you might've subtracted wrong @Britt101

OpenStudy (anonymous):

no you're right but its square root

OpenStudy (anonymous):

now you have to solve for x when its + and when its -

OpenStudy (comm.dan):

@galacticwavesXX Are you trying to say that she is right?

OpenStudy (anonymous):

im saying she hasn't solved for it completely

OpenStudy (comm.dan):

Ok. Just wanting to make sure. @galacticwavesXX

OpenStudy (anonymous):

no that as far as im suppose to go. I don't have to solve for x.

OpenStudy (anonymous):

she has \[\frac{ -9\pm \sqrt{69} }{ 2 }\]

OpenStudy (anonymous):

now she has to solve for x which shouldn't be too hard

OpenStudy (anonymous):

yeah that what i chose for my answer it dose not ask to go and find x.

OpenStudy (comm.dan):

@Britt101 Just solve for both of the x's and you will find it easy after you do some practice.

OpenStudy (anonymous):

oh ok just making sure!

OpenStudy (anonymous):

thanks you guys :) I really glad ya'll both helped :) Thanks.

OpenStudy (anonymous):

you welcome anytime!

OpenStudy (comm.dan):

No problem. Do you have any more questions at this time? @Britt101

OpenStudy (anonymous):

@galacticwavesXX What are the solutions of 4x2 – 3x – 5 = 0? is it \[\frac{3 \pm \sqrt{89}} {8} \]

OpenStudy (anonymous):

let me do it first see if you got it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

yes it is correct

OpenStudy (anonymous):

you got it!!:)

OpenStudy (comm.dan):

Nice job. Do u have any more regular or clarification questions? @Britt101

OpenStudy (anonymous):

yeah Solve. 3x2 = 147 x=7 or x=7 & x=-7? I think it's x=7&x=-7